OpenStudy (anonymous):

The Integral from 2 to sqrt2 of 1/(t^3sqrt(t^2-1))dt

OpenStudy (anonymous):

\[\int\limits_{2}^{\sqrt{2}}\frac{ 1 }{ t^3\sqrt{t^2-1} }dt\] ?

OpenStudy (anonymous):

Just the integral of 2 is on top.

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

reverse 2 and sqrt(2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok well it took a while but : \[t=coshx\] \[dt=sinhxdx\] then your integral become : \[\int\limits_{\sqrt2}^{2} \frac{ 1 }{ \cosh^3x }dx\] now using integration by parts \[\int\limits_{\sqrt2}^{2} \frac{ 1 }{ coshx }\frac{ 1 }{ \cosh^2x }dx\] \[v=\frac{ 1 }{ coshx } , u'=\frac{ 1 }{ \cosh^2x } \]

OpenStudy (anonymous):

\[v'=\frac{ -sinhx }{ \cosh^2x } , u=tanhx\] so we get by integration by parts : \[\int\limits_{}^{}\frac{ 1 }{ \cosh^3x }=\frac{ tanhx }{ coshx } + \int\limits_{}^{}\frac{ \sinh^2x }{ \cosh^3x }\] we can write it as : \[\int\limits_{}^{}\frac{ 1 }{ \cosh^3x }=\frac{ tanhx }{ coshx } + \int\limits_{}^{}\frac{ \cosh^2x-1 }{ \cosh^3x }\] then we have: \[2\int\limits_{}^{}\frac{ 1 }{ \cosh^3x }=\frac{ tanhx }{ coshx } + \int\limits_{}^{}\frac{ 1 }{ coshx }\] we divide by 2 and in order to solve 1/coshx i do : \[\int\limits_{}^{}\frac{ 1 }{ \cosh^3x }=\frac{ tanhx }{ 2coshx } + \frac{ 1 }{ 2 } \int\limits_{}^{}\frac{ 1 }{ coshx }\frac{ coshx }{ coshx }\] the n we get: \[\int\limits_{}^{}\frac{ 1 }{ \cosh^3x }=\frac{ tanhx }{ 2coshx } + \frac{ 1 }{ 2 } \int\limits_{}^{}\frac{ coshx }{ 1+\sinh^2x }\] in order to solve

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ coshx }{ 1+\sinh^2x }\] i use another sub p = sinhx so dp=coshxdx then this integral is easy and i get arctan(p) so we have \[\int\limits_{\sqrt2}^{2}\frac{ 1 }{ \cosh^3x } = \frac{ tanhx }{ 2coshx } + \frac{ 1 }{ 2 }\arctan(sinhx)\] now you can answer remembering coshx = t